# Where is the Super-Virial Gas? The Supply from hot inflows

### v1

@@ Abstract
In an effort to understand the presence of super-virial gas detected in the Milky Way, we present our findings from isolated galaxy simulations of Milky Way-like systems using GIZMO with the FIRE-2 (Feedback In Realistic Environments) stellar feedback model. It unveils the presence of a significant super-virial temperature ( K) gas component within 20 kpc from the galactic center. This super-virial gas has a mass of and is found close to the disk, where typical gas densities are . We find that some of the virial gas (K) forms a rotating hot inflow, where gravitational energy is converted to heat mainly via compressive heating. This process causes gas infalling close to the rotation axis to reach super-virial temperatures just before cooling and joining the disk. Stellar feedback heating accounts for less than 1% of the super-virial gas, indicating its minimal influence despite expectations. Even in scenarios with no stellar feedback effects considered, abundant super-virial gas persists, highlighting the dominance of alternative heating mechanisms. We also show that cosmic rays do not have a significant effect on heating the gas to a super-virial temperature. Our study illuminates the intricate dynamics of hot virial and super-virial gas surrounding Milky Way-like galaxies, emphasizing the prominent role of infall-driven and compressive heating processes in shaping thermal evolution.
UTF8mj
@@ 1 Introduction
Galaxies are the building blocks of our universe and how galaxies form and evolve is an active area of research. A spiral galaxy has two main components, a star-forming disk and a diffuse gaseous halo surrounding the disk. This diffuse gaseous halo, known as Circumgalactic medium (CGM) (Tumlinson et al., 2017; Faucher-Giguère & Oh, 2023), plays a crucial part in galaxy evolution. The CGM is the habitat for large-scale gas flows that fundamentally regulate the evolution of the galaxy by providing fresh and recycled gaseous fuel for star formation and regulating the galaxy’s interactions with other galaxies. A comprehensive understanding of the CGM stands to illuminate key unresolved questions pertaining to the processes driving galaxy formation and evolution.
@@
Unfortunately, the CGM is very hard to detect in emission due to its very low density. However, from absorption lines in the spectra of background quasars, we can infer about the physical properties of the CGM. When light from a background quasar passes through the CGM of a galaxy, it suffers absorption from the metals and ions in the CGM. Absorption studies made by COS spectrograph onboard HST along with XMM-Newton and Chandra has enabled us to detect this multiphase diffuse gas around Milky Way (MW) mass galaxy in UV [e.g. Tumlinson et al. (2011); Werk et al. (2016)] and X-ray [e.g. Gupta et al. (2012); Fang et al. (2015); Das et al. (2021); Mathur (2022)] respectively. In the X-ray band, the MW CGM has been detected in X-ray emission as well [e.g. Snowden et al. (2000); Henley & Shelton (2013); Nakashima et al. (2018).]
@@
Untill 2019, These observations led us to a picture where the CGM is multi-phase in nature and has three main components; virial temperature warm-hot gas (T K) , warm gas (T K) and cool gas (T K). However this picture got modified with the recent discovery of the super-virial component (TK; Das et al. (2019b)) in the MW CGM using absorption lines from XMM-Newton observation towards the Blazar IES 1553+113. Since then, there have been multiple detections of this hotter super-virial gas in both emission and absorption (Das et al., 2019a; Gupta et al., 2023; Bhattacharyya et al., 2023; Lara-DI et al., 2023; McClain et al., 2024). In emission, Das et al. (2019a) detected the super-virial component towards the same sight-line of Blazar IES 1553+113 using XMM-Newton, and Gupta et al. (2023) found this component along with virial component along four sightlines from Suzaku and Chandra. Later, Bhattacharyya et al. (2023) also detected this super-virial gas in the MW CGM in emission towards the Blazar Mrk 421 and 5 other sightlines near the Blazar field. Even in a recent study by McClain et al. (2024) detected SiXIV K and SXVI K, associated with a gas phase with super-virial temperature, in the CGM of the MW, with the use of stacked X-ray spectral observations towards multiple of different extragalactic sightlines.
@@
This super-virial component has not only been detected towards some particular sightlines, but all-sky surveys from Halo-Sat and eROSITA have also found this component abundantly in all over the sky. Bluem et al. (2022) utilized data from the HaloSat all-sky survey to delve into the emissions originating from the MW’s halo and detected the supervirial gas. Ponti et al. (2023), in their study of the MW halo within the eFEDDS field employing eROSITA, also detected this component.
@@
However, the gas at these temperatures defies the predictions of galaxy formation simulations (e.g., Shen et al. (2013)). Its discovery begs the questions like: where is this gas situated; in the ISM, inner CGM or outer CGM? How this gas is formed; through feedback or some other heating mechanisms like cosmic ray (CR) heating? What is the nature of this gas; is it as diffuse as the volume filling virial temperature gas or is it more or less dense?
@@
In this paper, we will address above questions in light of an idealized simulation. We investigate a suite of simulations of MW-type hosts where we focus on these super-virial gas components. The paper is structured as follows. In Section 2, we discuss the methodology of our simulation, where we describe the initial conditions of our simulation setup (in Section 2.1) along with our definition of super-virial CGM gas for our analysis (in Section 2.2). In Section 3, we demonstrate our results from our analysis of the simulations, where we discuss the location (Section 3.1), density (in Section 3.2), amount (Section 3.3), and origin (Section 3.4) of the super-virial gas in the CGM. We also discuss if feedback and cosmic rays heating have any effect on the origin of this gas in Section 3.1 and 3.3. We discuss the limitations of our work and planned future work in Section 4. Finally, we summarize our results and discuss future work in Section 5.
Symbol | Resolution | Feedback | Cosmic Ray |
m12_lr | low | Yes | No |
m12_hr | high | Yes | No |
m12_lr_nofb | low | No | No |
m12_lr_CR_low | low | Yes | Yes (low) |
m12_lr_CR_high | low | Yes | Yes (high) |
m12_lr_CF (cooling flow) | low | Yes | No |
@@ 2 Methodology
In all simulations except for the runs with no feedback, we incorporate the FIRE-2 model, an implementation of the Feedback In Realistic Environments (FIRE) approach. This encompasses treatments of the interstellar medium (ISM), star formation, and stellar feedback, the details of which are given in Hopkins et al. (2018a, b) along with extensive numerical tests. Our simulations cover a broad temperature range from to K, accounting for various cooling mechanisms including photoelectric and photoionization heating, collisional, Compton, fine structure, recombination, atomic, and molecular cooling.
@@
The model for cosmic ray (CR) treatment includes streaming, which occurs at the local Alfvén speed or sound speed, whichever is larger, incorporating an appropriate streaming loss term that thermalizes energy, as described in Uhlig et al. (2012). Diffusion is modeled with a fixed diffusivity , alongside adiabatic energy exchange between gas and CR pressure, and includes hadronic and Coulomb losses (following Guo & Oh (2008)). We consider a single energy bin for GeV CRs, which dominate the pressure, and treat them in the ultra-relativistic limit. Both streaming and diffusion are fully anisotropic along magnetic field lines. CRs are injected by supernovae (SNe), with 10% of each SNe energy being transferred into CRs, consistent with studies like Pfrommer et al. (2017a, b). For details on cosmic ray physics is demonstrated in Su et al. (2019a); Chan et al. (2019b).
@@ 2.1 Initial conditions
The initial setup closely follows the detailed specifications outlined in Su et al. (2019b); Roy et al. (2024). To further stabilize the host CGM, the simulation region is expanded to 3 times the viral radius, and the simulations are run adiabatically (no cooling or star formation) for 4.5 Gyr for relaxing any initial transients before the satellites are placed into the CGM. The simulation properties are summarized in table 1 of Roy et al. (2024). In this paper, we only focus on the m12 halo which is corresponding to the MW-like halo of mass .
We consider only the runs with 20 satellites of DM halo mass (20xm09). We vary different physical properties to test the impact of the stellar feedback and cosmic rays. We summarize all of our runs in Table 1.
We initialize the dark matter (DM) halo, bulge, black hole, and gas+stellar disk of m12 by following Springel & White (1999); Springel (2000). The initial metallicity goes down from solar () to with radius as . Magnetic fields are initialized azimuthal with . For our cooling flow run, we have the same initial condition but with a cooling flow solution with cooling flow of roughly 5 M yr-1, a sonic radius of 1 kpc and a circularized radius of 10 kpc (following Stern et al., 2024). We initialize CRs in the high CR runs such that the CR energy density is in equipartition with thermal energy density. In this case, we decrease the temperature by half, keeping the density same, and put the rest of the energy in CR. So the pressure gradient is maintained. In low CR runs, CRs are initialized such that the CR energy density is in equipartition with magnetic energy density, which is order of magnitude lower than the thermal energy density.
We initialize the satellites similarly as the host described above, though without a CGM gas halo. The properties are also summarized in table 1 of Roy et al. (2024). More details of initial set up is described in Methodology section of Roy et al. (2024).
@@ 2.1.1 Gas temperature of the initial condition
All the runs have similar initial setup except for the high CR run. We check that our initial setups for different runs do not contain any super-virial gas. We confirm that the initial temperature is not more than K in all the runs. In the high CR run, the initial temperature profile is even lower for maintaining hydro-static pressure balance, not crossing K. The variation in the initial condition will be useful to test whether the super-virial gas can be found even with a cooler halo, and test whether the results are sensitive to the initial set-up. We also confirms that the initial density ranges similarly in all the cases, with virial gas having density .
@@ 2.2 Defining Super-virial Gas
Observationally, the gas with temperature K has been identified as super-virial phase. For this analysis, we use temperature cut of K for super-virial gas. Therefore, we consider gas with temperature above this threshold to be called as ‘super-virial’.
@@ 3.1 Where is the super-virial gas?
This section delves into the investigation of the distribution of supervirial gas within the simulation box. We present the 2-d PDFs of temperature of the gas weighted by its mass with the variation of its position and density, in the left and right panels of Figure 1 respectively. The 2d-PDFs denote the four different runs at 1 Gyr snapshot of the simulation for no CR (1st row), no CR and no Feedback (2nd row), low CR (3rd row), and high CR (bottom panel) cases.
In all the 2-d distribution plots, we did a radial cut of 6 times of the gas scale radius region around the satellites to exclude the ISM of the satellites, as we choose this radius of the satellite to be the radius beyond which the gravitational pull from the satellites is negligible. Note, however, that we do not exclude satellite particles which get stripped from the satellites. The horizontal line in Figure 1 denotes the super-virial temperature cut of K; gas with temperature above this threshold is called ‘super-virial’.
In the right bar of each 2-d histogram plot in Figure 1, we show the 1-d histograms of mass-weighted temperature at two different positions (left panels) and density (right panels) denoted by the similarly colored vertical lines in the 2-d histograms. For example, in the left panel of Figure 1, there is a red vertical line at 20 kpc and the corresponding histogram in the right bar, showing the temperature distribution at 20 kpc. We see that the distribution not extends above K. On the other hand, the temperature distribution shown by the black histogram (corresponding to 6.5 kpc) extends far beyond K. Therefore, we conclude that the majority of the super-virial gas is concentrated within a 20 kpc radius across all of the simulation runs.
We also checked if we can see this gas at an earlier snapshot at t=0.5 Gyr and we found it for all of cases except for the high CR one, where the super-virial gas is not there at 0.5 Gyr, but appears at 1 Gyr. This is mainly due to the fact that the high CR runs have cooler initial halo, therefore taking longer time to get heated.
In summary, we investigate the distribution of supervirial gas within the simulation box, initially examining its temperature and density distributions. Excluding regions around satellites, the analysis reveals that the majority of super-virial gas, exceeding K, is concentrated within a 20 kpc radius across all runs. Notably, the high CR runs exhibit delayed appearance of super-virial gas at 1 Gyr due to cooler initial halo conditions, highlighting sensitivity to initial setup and heating dynamics.
@@ 3.2 How dense is the super-virial gas?
In this section, we scrutinize the density distribution of the super-virial gas. In the right panel of Figure 1, we present 2D histograms illustrating the relationship between temperature and density, weighted by mass. It shows the temperature evolution of 2-d probability distribution function (PDF) of gas mass as a function of the temperature and density of the gas at t=1 Gyr for no CR (1st row), no CR and no Feedback (2nd row), low CR (3rd row), and high CR (bottom panel) cases. In these plots, we excluded the region around the satellites like before. Notably, in all the cases, the density of the super-virial gas is approximately as the gas is closer to the disk, as shown in left panel of Figure 1.
@@ 3.3 How much super-virial gas is there in the galaxy?
Next, we quantify the mass of super-virial gas present in the galaxy. Given that the bulk of the super-virial gas resides within a 20 kpc radius, we depict the temporal evolution of super-virial gas masses within 20 kpc and 10 kpc using solid and dashed lines, respectively, in Figure 2. Here, we do not exclude the region around the satellites, as we would like to check if satellites also contributes to this gas phase. However, as the satellites are in circular orbits between 100kpc and 150kpc, the regions around them are automatically cut out while we calculate super-virial gas mass inside 20 kpc and 10 kpc. Key insights from this analysis include:
@@
Mass: The mass of super-virial gas fluctuates within the range of throughout the entire 1.5 Gyr simulation period within 10 kpc, with 2 times larger mass within 20kpc. CGM gas mass being , the supervirial gas is only 0.1-1% of the total CGM gas, hence sub-dominant in the baryonic budget and does not contribute to the missing baryon problem.
@@
Effect of Feedback: The presence of the super-virial gas could have been a direct result of heating by stellar feedback. Therefore, if we don’t include feedback in our simulations, we might expect no (or significantly smaller amount of) super-virial gas. Surprisingly, however, the simulations with no feedback display a higher and similar amount of mass of super-virial gas among all the runs. This shows that in our simulations, the stellar feedback is not predominantly responsible for generating the super-virial gas.
Effect of Cosmic Rays: CR heating by streaming heating or Coulomb heating can be one of the heating source of the super-virial gas. It can also drive outflow that can heat the gas. We investigate this with two runs with CRs, one with low CR and one with high CR. Minimal disparities exist between CR and no CR runs. The runs incorporating CRs exhibit slightly more pronounced mass fluctuations compared to runs without CRs, with the high CR run showing an even larger fluctuations. Hence, in our simulations, the CR is not predominantly responsible for heating the gas to super-virial phase.
Effect of Resolution: We have conducted higher-resolution simulations (10 times higher resolution compared to our low resolution run), represented by dashed lines in Figure 2. Our anticipation was that these high-resolution (hr) runs would capture denser,super-virial gas compared to the low-resolution (lr) counterparts. We observe that the super-virial gas mass in the hr runs qualitatively similar to that in the lr runs. This consistency suggests that our findings remain robust across varying levels of resolution.
@@
Effect of Other Satellite Runs: We also check the other satellite runs described in Roy et al. (2024), with m10 and m08 satellites to check whether the satellite distribution has a significant effect on our results (not shown in the figure). We find that there is no significant difference in the super-virial gas mass budget with the difference in satellite distribution.
@@
In summary, we find that the temporal evolution of super-virial gas mass within 20 and 10 kpc radii reveals fluctuations between , with a doubling of mass within the broader radius, constituting only 0.1-1% of the total CGM gas mass. Surprisingly, simulations without feedback display higher and more consistent super-virial gas masses, suggesting super-virial gas does not predominantly originate from feedback heating processes. Minimal disparities in super-virial gas mass fluctuations are observed between runs with and without cosmic rays, with higher CR runs showing slightly more pronounced fluctuations.

### v2

@@ Abstract
To understand the presence of the super-virial temperature gas detected in the Milky Way, we present our findings from isolated galaxy simulations of Milky Way-mass systems using GIZMO with the FIRE-2 (Feedback In Realistic Environments) stellar feedback model. It unveils the presence of a significant super-virial temperature () gas component within 20 kpc from the galactic center. We also find that 70–90% of the total super-virial gas is extra-planar, at kpc and kpc. This super-virial gas has a mass of with typical gas densities are . We find that some of the virial gas (K) forms a rotating hot inflow, where gravitational energy is converted to thermal energy mainly via compressive heating. This process causes gas falling close to the rotation axis to reach super-virial temperatures via a combination of compressive heating and shocks just before cooling and joining the disk. Stellar feedback heating accounts for less than 1% of the super-virial gas, indicating its minimal influence despite expectations. Even in scenarios with no stellar feedback effects considered, abundant super-virial gas persists, highlighting the dominance of alternative heating mechanisms. We also show that cosmic rays do not have a significant effect on heating the gas to a super-virial temperature. Our study illuminates the intricate dynamics of hot virial and super-virial gas surrounding Milky Way-mass galaxies, emphasizing the prominent role of infall-driven compressive and shock-heating processes in shaping thermal evolution.
@@ 1 Introduction
Galaxies are the building blocks of our universe, and how galaxies form and evolve is an active area of research. A spiral galaxy has two main components: a star-forming disk and a diffuse gaseous halo surrounding the disk. This diffuse gaseous halo, known as the Circumgalactic medium (CGM) (Tumlinson et al., 2017; Faucher-Giguère & Oh, 2023), plays a crucial part in galaxy evolution. The CGM is the habitat for large-scale gas flows that fundamentally play an important role in the evolution of the galaxy by providing fresh and recycled gaseous fuel for star formation and regulating the galaxy’s interactions with other galaxies. A comprehensive understanding of the CGM stands to illuminate key unresolved questions about the processes driving galaxy formation and evolution.
@@
Unfortunately, CGM is very hard to detect in emission due to its low density. However, from absorption lines in the spectra of background quasars, we can infer the physical properties of the CGM. When light from a background quasar passes through the CGM of a galaxy, it suffers absorption from the metals and ions in the CGM. Absorption studies made by COS spectrograph onboard HST along with XMM-Newton and Chandra have enabled us to detect this multiphase diffuse gas around Milky Way (MW)-like galaxy in UV [e.g. Tumlinson et al. (2011); Werk et al. (2016)] and X-ray [e.g. Gupta et al. (2012); Fang et al. (2015); Das et al. (2021); Mathur (2022)] respectively. The MW CGM has also been detected in X-ray emission [e.g., Snowden et al. (2000); Henley & Shelton (2013); Nakashima et al. (2018)].
@@
Until 2019, these observations led us to a picture where the CGM is multi-phase in nature and has three main components; virial temperature warm-hot gas (T K), warm gas (T K) and cool gas (T K). However, this picture got modified with the recent discovery of the super-virial component (TK; Das et al. (2019b)) in the MW CGM using absorption lines from XMM-Newton observation towards the Blazar IES 1553+113. Since then, there have been multiple detections of this hotter super-virial gas in both emission and absorption (Das et al., 2019a; Gupta et al., 2023; Bhattacharyya et al., 2023; Lara-DI et al., 2023; McClain et al., 2024). In emission, Das et al. (2019a) detected the super-virial component towards the same sight-line of Blazar IES 1553+113 using XMM-Newton, and Gupta et al. (2021) found this component along with the virial component toward four sightlines from Suzaku and Chandra. Later, Bhattacharyya et al. (2023) also detected this super-virial gas in the MW CGM in emission toward the Blazar Mrk 421 and 5 other sightlines near the Blazar field. Even a recent study by Lara-DI et al. (2023) detected SiXIV K and SXVI K, associated with a gas phase with super-virial temperature, in the CGM of the MW, with the use of stacked X-ray spectral observations towards multiple different extragalactic sightlines.
@@
The super-virial component has not only been detected towards some particular sightlines, but all-sky surveys from Suzaku, Halo-Sat, and eROSITA have also found this component abundantly all over the sky. Gupta et al. (2023) found this component abundantly while analyzing the X-ray emission of the shells of the eROSITA bubbles from Suzaku data. Bluem et al. (2022) utilized observations from the HaloSat all-sky survey to delve into the emissions from the MW’s halo and detected the super-virial gas. Ponti et al. (2023), in their study of the MW halo within the eFEDDS field employing eROSITA, also detected this component.
@@
This discovery begs questions like: where is this gas situated; in the ISM, inner CGM, or outer CGM? How is this gas formed: through feedback or some other heating mechanisms like cosmic ray (CR) heating? What is the nature of this gas: is it as diffuse as the volume-filling virial temperature gas, or is it more or less dense?
@@
In this paper, we will address the above questions in light of an idealized simulation. We investigate a suite of simulations of MW-mass host galaxy, where we focus on this super-virial gas component. The paper is structured as follows. In Section 2, we discuss the methodology of our simulation, where we describe the initial conditions of our simulation setup (in Section 2.1) along with our definition of super-virial CGM gas for our analysis (in Section 2.2). In Section 3, we demonstrate our results from our analysis of the simulations, where we discuss the location (Section 3.1), density (in Section 3.2), amount (Section 3.3), and origin (Section 3.4) of the super-virial gas in the CGM. We also discuss if feedback and cosmic rays heating have any effect on the origin of this gas in Section 3.1 and 3.3. We discuss the limitations of our work and planned future work in Section 4. Finally, we summarize our results and discuss future work in Section 5.
@@ 2 Methodology
In all simulations except for the runs with no feedback, we incorporate the FIRE-2 model, an implementation of the Feedback In Realistic Environments (FIRE) approach. This encompasses treatments of the interstellar medium (ISM), star formation, and stellar feedback, the details of which are given in Hopkins et al. (2018a, b) along with extensive numerical tests. Our simulations cover a broad temperature range from to K, accounting for various cooling mechanisms, including photoelectric and photo-ionization heating, collisional, Compton, fine structure, recombination, atomic, and molecular cooling.
@@
The model of cosmic ray (CR) treatment includes streaming, which occurs at the local Alfvén speed or sound speed, whichever is larger, incorporating an appropriate streaming loss term that thermalizes energy, as described in Uhlig et al. (2012). Diffusion is modeled with a fixed diffusivity alongside adiabatic energy exchange between gas and CR pressure and includes hadronic and Coulomb losses (following Guo & Oh (2008)). We consider a single energy bin for GeV CRs, which dominate the pressure, and treat them in the ultra-relativistic limit. Both streaming and diffusion are fully anisotropic along magnetic field lines. CRs are injected by supernovae (SNe), with 10% of each SNe’s energy being transferred into CRs, consistent with studies like Pfrommer et al. (2017a, b). Details on cosmic ray physics are demonstrated in Su et al. (2019); Chan et al. (2019b).
@@ 2.1 Initial conditions
The initial setup follows the detailed specifications outlined in Su et al. (2019); Roy et al. (2024). To further stabilize the host CGM, the simulation region is expanded to 3 times the viral radius, and the simulations are run adiabatically (no cooling or star formation) for 4.5 Gyr to relax any initial transients before the satellites are placed into the CGM. The simulation properties are summarized in Table 1 of Roy et al. (2024). In this paper, we only focus on the m12 halo which is corresponding to the MW-mass halo of mass .
We consider only the runs with 20 satellites of DM halo mass (20xm09).
We vary different physical properties to test the impact of stellar feedback and cosmic rays. Our runs are the following :
m12_lr: low resolution () and without CR
m12_hr: high resolution () and without CR
m12_lr_nofb: low resolution (), without CR and feedback. It is important to note that a galaxy without feedback can not represent a realistic system; it is therefore not intended to match observed galaxies, but rather serves as a baseline to isolate and highlight the role of feedback.
m12_lr_CR_low: low resolution () and with low CR
m12_lr_CR_high: low resolution () and with high CR
We initialize the dark matter (DM) halo, bulge, black hole, and gas+stellar disk of m12 by following Springel & White (1999); Springel (2000). The initial metallicity goes down from solar () to with radius as . Magnetic fields are initialized azimuthal with . More details of the initial setup are described in the Methodology section of Roy et al. (2024). We initialize CRs in the high CR runs such that the CR energy density is in equipartition with the thermal energy density. In this case, we decrease the temperature by half, keeping the density the same, and put the rest of the energy in CR. In low CR runs, CRs are initialized such that the CR energy density is in equipartition with magnetic energy density, which is order of magnitude lower than the thermal energy density.
We initialize the satellites similarly, like the host described above, though without a CGM gas halo. The satellites are initialized orbiting randomly between 100 kpc to 150 kpc of the host galaxy in a circular orbit with the circular velocity of the host galaxy at that radius. The properties are also summarized in Table 1 of Roy et al. (2024). The satellite configuration was originally designed to see satellite contribution to CGM properties, and we find that they have relatively little contribution in producing the super-virial phase. Therefore, we focus mainly on the host gas for this work.
@@ 2.1.1 Gas temperature of the initial condition
All the runs have similar initial setups, except for the high CR run. We check that our initial setups for different runs do not contain any super-virial gas. We confirm that the initial temperature is not more than K in all the runs. In the high CR run, the initial temperature profile is even lower for maintaining hydrostatic pressure balance, not crossing K. The variation in the initial condition will be useful to test whether the super-virial gas can be found even with a cooler halo, and test whether the results are sensitive to the initial setup. We also confirm that the initial density ranges similarly in all the cases, with virial gas having density .
@@ 2.2 Defining Super-virial Gas
Observationally, the gas with a temperature greater than the virial temperature has been identified as the super-virial phase. For this analysis, we use the temperature cut of K for the super-virial temperature gas. Thus, we consider gas with a temperature above this threshold as the ‘super-virial’ gas.
@@ 3.1 Where is the super-virial gas?
This section delves into the investigation of the distribution of super-virial gas within the simulation box. We present the temperature distribution of gas in the cylindrical radius (Rcyl) and galactic height (z) plane, in the top row of Figure 1. The distribution denotes the three different runs at a 1 Gyr snapshot of the simulation for no CR (1st column), no CR and no Feedback (2nd column), and the low CR (3rd column) cases. In all the 2-D distribution plots, we did a radial cut of 6 times the gas scale radius region around the satellites to exclude the ISM of the satellites, as we chose this radius of the satellite to be the radius beyond which the gravitational pull from the satellites is negligible. Note, however, that we do not exclude satellite particles that can be stripped from the satellites.
The top panel shows that the super-virial phase is in the extra-planar region in all the cases, at and kpc. Note that 6 kpc is the gas disc exponential scale length of the galaxy, and 1 kpc is the stellar bulge Hernquist-profile scale length. Thus, we see that the super-virial gas mainly extends up to the disc gas scale length but beyond the stellar bulge scale length for a galaxy. However, the different physical processes lead to different extra-planar geometries. With no feedback, the super-virial gas is in a flat geometry and distributed more in Rcyl. With stellar feedback, the gas is more concentrated at inner Rcyl and inner z. With CR feedback, the gas is more extended in z.
@@ 3.2 How dense is the super-virial gas?
In this section, we scrutinize the density distribution of the super-virial gas. In the bottom panel of Figure 1, we show the temperature density 2-d distribution weighted by mass. The horizontal line in the bottom panel of Figure 1 denotes the super-virial temperature cut of K; gas with the temperature above this threshold is called ‘super-virial’. The temperature-density phase-space of the gas shows that the density of the super-virial gas is roughly in between cm-3. This density is higher than the virial volume filling gas density as the gas is closer to the disk, as shown in the top panel of Figure 1. Note that the no-feedback run has a pillar of lower-density gas in the super-virial phase, arising from the more distributed geometry mentioned in the above section.
@@ 3.3 How much super-virial gas is there in the galaxy?
Next, we quantify the mass of super-virial gas present in the galaxy. We plot the total mass of the super-virial gas in the top-left panel of Figure 2 for all the runs. Given that the bulk of the super-virial gas resides within R kpc and kpc, we plot the fraction of total super-virial gas present within this region in the top-right panel of Figure 2. Here, we do not exclude the region around the satellites, as we would like to check if satellites also contribute to this gas phase. As the satellites are in circular orbits between 100 kpc and 150 kpc, the regions around them are automatically cut out while we calculate the super-virial gas mass inside 20 kpc and 10 kpc. Key insights from this analysis include:
@@
Mass: The mass of total super-virial gas fluctuates within the range of throughout the entire 1.5 Gyr simulation period. CGM gas mass being , the super-virial gas is only 0.1-1% of the total CGM gas, hence sub-dominant in the baryonic budget and does not contribute to the missing baryon problem. The top-right panel of Figure 2 shows that 70–90% of total super-virial gas is within kpc and kpc, again supporting the extra-planar geometry.
@@
Effect of Feedback: The presence of the super-virial gas could have been a direct result of heating by stellar feedback. Therefore, if we do not include feedback in our simulations, we might expect no (or a significantly smaller amount of) super-virial gas. Surprisingly, however, the simulations with no feedback display a higher amount of mass of super-virial gas among all the runs. It also indicates that in our simulations, the stellar feedback is not predominantly responsible for generating the super-virial gas.
Effect of Cosmic Rays: CR heating by streaming heating or Coulomb heating can be one of the heating sources of the super-virial gas. CR can also drive outflows that can heat the gas. We investigate these effects with two runs with CRs, one with low CR and one with high CR. The mass of super-virial gas exhibits greater temporal fluctuations with increasing cosmic ray (CR) feedback, although the overall magnitude of the mass budget remains qualitatively similar. This suggests that while CR feedback induces significant variability in the super-virial gas content, it does not directly contribute to its origin.
Effect of Resolution: We have conducted higher-resolution simulations (10 times higher resolution compared to our low-resolution run), represented by light-shaded blue line in Figure 2. We anticipate that these high-resolution (hr) runs would capture denser, super-virial gas than the low-resolution (lr) counterparts. In the top-left panel of Figure 2, we observe that the total super-virial gas mass in the hr runs qualitatively similar to that in the lr runs. The top-right panel of Figure 2 shows the amount of gas within a 15 kpc radius and 1–6 kpc height. Here, we see a 20% difference between the high- and low-resolution simulations. The remaining 20% super-virial gas in the high-resolution run is present at kpc. This is because the denser and hotter structures at higher galactic heights are resolved by the high-resolution run. While these results demonstrate improved capture of vertical structure in the high-resolution run, we are currently limited by computational constraints and can not explore even higher resolutions. Overcoming this limitation is an important direction for future work.
@@
Effect of Other Satellite Runs: We also check the other satellite runs described in Roy et al. (2024), with m10 and m08 satellites, to check whether the satellite distribution has a significant effect on our results (not shown in the figure). We find that there is no significant difference in the super-virial gas mass budget with the difference in satellite distribution. However, we discuss how our result can depend on the satellites’ initial conditions in the discussion section.
@@
In summary, we find that the total super-virial gas mass lies in the range as the galaxy evolves. The simulations without feedback display higher super-virial gas masses, suggesting super-virial gas does not predominantly originate from feedback heating processes. Minimal disparities in the super-virial gas mass fluctuations are observed between runs with and without cosmic rays, with higher CR runs showing slightly more pronounced fluctuations. of the total super-virial gas lies within Rkpc and kpc, suggesting an extra-planar geometry extending up to kpc scale height.